OpenStudy (anonymous):

absolute convergence , converge or diverge: n^2/2^n. sumation going from n=1 until infinity

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} {\frac{n^2}{2^n}}\]

OpenStudy (anonymous):

Check it with the following test : \[\lim_{n\to \infty} {\frac{u^{n+1}}{u^n}}\]

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

sr, it is \[\lim_{n\to \infty} {\frac{u_{n+1}}{u_n}}\]

OpenStudy (anonymous):

with \[u_n=\frac{n^2}{2^n}\] here

OpenStudy (anonymous):

wats sr?

OpenStudy (anonymous):

i dont understand the formula ur using

OpenStudy (anonymous):

this is d'Alembert test

OpenStudy (anonymous):

if this limit < 1, the series is converge

OpenStudy (anonymous):

neverheard of that...aight

OpenStudy (anonymous):

what is the test you studied?

OpenStudy (anonymous):

we did power, ratio?

OpenStudy (anonymous):

It seems to be a ratio test :-? may be differ name, display ratio test here

OpenStudy (anonymous):

given sumation A sub n. let the ratio= lim n goes to infinity of absolute val of A sub n +1 / A sub n then if ratio<1 series Coverges Absloutely ratio>1, diverges ratio=1, try diff test

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

so how do u write it out?

OpenStudy (anonymous):

\[\lim_{n\to \infty} {\frac{\frac{(n+1)^2}{2^{n+1}}}{\frac{n^2}{2^n}}}=\lim_{n\to \infty} {\frac{(n+1)^2}{2n^2}}=\frac{1}{2}<1\]

OpenStudy (anonymous):

howo does thatcancel out?

OpenStudy (anonymous):

I don't understand what you say, but if this limit <1, you can deduce this series converges. Moreover, it is absolute converge because \[\frac{n^2}{2^n}>0\]

OpenStudy (anonymous):

how do you get that second part of the limit?

OpenStudy (anonymous):

\[\lim_{n\to \infty} {\frac{(n+1)^2}{2n^2}}=\frac{1}{2}\]???

OpenStudy (anonymous):